# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v1

@@ IV-B Infeasibility Detection
The vector $\mathbf{v}$ can be written as $\mathbf{v}=A^{T}\bm{\lambda},\;\bm{\lambda}\in^{n_{C}}$. For all $\bm{\xi}\in\mathcal{A}$, $A\bm{\xi}=\mathbf{b}$, and therefore $\mathbf{v}^{T}\bm{\xi}=\bm{\lambda}^{T}\mathbf{b}$ $\forall\bm{\xi}\in\mathcal{A}$ and $\min_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})=\max_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})=\mathbf{v}^{T}\tilde{\bm{\xi}}$. Defining $[c]=\mathbf{v}^{T}[\bm{\beta}]=[\underline{c},\overline{c}]$, $\min_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})=\underline{c}$ and $\max_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})=\overline{c}$. If $\mathbf{v}^{T}\tilde{\bm{\xi}}\notin[c]$, then $\min_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})>\max_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})$ or $\min_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})>\max_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})$, and by the separating hyperplane theorem $\mathcal{A}\cap\mathcal{B}=\emptyset$ [33], so (22) is infeasible. ∎
@@
If (22) is infeasible and $Z^{T}\tilde{P}Z$ is positive definite where $Z$ is a basis for the null space of $A$, then the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}\rightarrow\mathbf{v}$ for $k\geq k^{\prime}$ with $k^{\prime}$ sufficiently large, where $\mathbf{v}$ is an infeasibility certificate for problem (22).
@@
From [31, Thm. 1], if (22) is infeasible and $Z^{T}\tilde{P}Z>0$, then for $k>k^{\prime}$, $\bm{\zeta}_{k}-\bm{\xi}_{k}=\bm{\lambda}^{o}\in\text{span}{(A^{T})}$ where $\bm{\lambda}^{o}=\bm{\zeta}^{o}-\bm{\xi}^{o}$ with $\bm{\xi}^{o}$ and $\bm{\zeta}^{o}$ the solutions to $\operatorname*{\mathrm{argmin}}_{\bm{\xi},\bm{\zeta}}||\bm{\zeta}-\bm{\xi}||^{2}$ s.t. $\bm{\xi}\in\mathcal{A}$ and $\bm{\zeta}\in\mathcal{B}$. Since (22) is infeasible, $||\bm{\zeta}^{o}-\bm{\xi}^{o}||^{2}=(\bm{\zeta}^{o}-\bm{\xi}^{o})^{T}(\bm{\zeta}^{o}-\bm{\xi}^{o})>0$. Because $\bm{\xi}^{o}$ and $\bm{\zeta}^{o}$ minimize $||\bm{\zeta}^{o}-\bm{\xi}^{o}||^{2}$ for $\bm{\xi}\in\mathcal{A}$ and $\bm{\zeta}\in\mathcal{B}$, $(\bm{\zeta}^{o}-\bm{\xi}^{o})^{T}(\bm{\zeta}-\bm{\xi})\geq(\bm{\zeta}^{o}-\bm{\xi}^{o})^{T}(\bm{\zeta}^{o}-\bm{\xi}^{o})>0$ $\forall\bm{\xi}\in\mathcal{A}$ and $\bm{\zeta}\in\mathcal{B}$. Thus, $(\bm{\lambda}^{o})^{T}\bm{\xi}<(\bm{\lambda}^{o})^{T}\bm{\zeta}$ $\forall\bm{\xi}\in\mathcal{A}$ and $\bm{\zeta}\in\mathcal{B}$ and by Theorem 1, $\mathbf{v}=\bm{\lambda}^{o}$ constitutes an infeasibility certificate. ∎
@@
IV-C Problem Scaling and Pre-conditioning
IV-D Interpretation of ADMM Iterations
@@
IV-E ADMM Algorithm for Constrained Zonotopes
@@
2:  $k\leftarrow 0$
3:  $(\bm{\xi}_{k},\bm{\zeta}_{k},\mathbf{u}_{k})\leftarrow(\mathbf{0},\mathbf{0},\mathbf{0})$
@@
15:     converged $\leftarrow(||\mathbf{r}_{p}||_{2}<\sqrt{n}\epsilon_{p})$ and $(||\mathbf{r}_{d}||_{2}<\sqrt{n}\epsilon_{d})$
@@ V-A Model Predictive Control
1:  $\mathcal{Z}\leftarrow\{\mathbf{x}_{0}\}$
2:  $P\leftarrow Q$
3:  $\mathbf{q}\leftarrow\mathbf{0}$
@@
7:        $P\leftarrow\text{blkdiag}\left(\begin{bmatrix}P&R&Q_{N}\end{bmatrix}\right)$
@@
9:        $P\leftarrow\text{blkdiag}\left(\begin{bmatrix}P&R&Q\end{bmatrix}\right)$
@@
[…]
@@ V-B Moving Horizon Estimation and Set-Valued State Estimation
$\mathbf{x}_{k+1}=A\mathbf{x}_{k}+B\mathbf{u}_{k}+\mathbf{w}_{k}\;,\;\mathbf{y}_{k}=\mathbf{x}_{k}+\mathbf{v}_{k}\;,$ (27a)
[…]
subject to bounded process and measurement noise $\mathbf{w}_{k}$ and $\mathbf{v}_{k}$, respectively. The state domain set $\mathcal{S}$ is discussed in Sec. III. Constrained zonotopes have been used to find a bounding set $\mathcal{X}_{k}$ such that $\mathbf{x}_{k}\in\mathcal{X}_{k}$ [9]. This related problem is known as set-valued state estimation (SVSE). While SVSE algorithms produce a set that is guaranteed to contain the state of the system, they do not in general produce a best state estimate within that set. The following presents an MHE implementation based on reachability analysis with constrained zonotopes that produces both a bounding set of system states and a best estimate of the state within that set.
@@
$\mathcal{X}_{k+1}=\Big{(}(A\mathcal{X}_{k}\oplus B\mathbf{u}_{k}\oplus\mathcal{W})\cap(\mathbf{y}_{k+1}\oplus(-\mathcal{V}))\Big{)}\cap\mathcal{S}\;,$ (28)
@@
$\mathcal{X}_{k+1}=\begin{bmatrix}0&0&I\end{bmatrix}\Big{(}(\mathcal{X}_{k}\times\mathcal{W}\times(\mathcal{S}\cap(\mathbf{y}_{k+1}\oplus(-\mathcal{V}))))\\ \cap_{\begin{bmatrix}A&I&-I\end{bmatrix}}\{-B\mathbf{u}_{k}\}\Big{)}\;.$ (29)
@@
$\mathbf{x}_{k+1}=A\mathbf{x}_{k}+B\mathbf{u}_{k}+\mathbf{w}_{k}\;,\;\mathbf{y}_{k}=\mathbf{x}_{k}+\mathbf{v}_{k}\;,$ (30b)
@@
The process and measurement noise covariance matrices are $Q$ and $R$, respectively. The covariance of the state estimate at time step $k=-N$ is $P_{-N}$, and the prior state estimate for $k=-N$ is $\hat{\mathbf{x}}_{-N}$. The set of possible states at time step $k=-N$ is $\mathcal{X}_{-N}$, which is updated recursively between MHE iterations by applying (29).
@@
1:  $\mathcal{Z}\leftarrow\mathcal{X}_{-N}$
2:  $P\leftarrow P^{-1}_{-N}$
3:  $\mathbf{q}=-P_{-N}^{-1}\hat{\mathbf{x}}_{-N}$
@@
6:     $P\leftarrow\text{blkdiag}\left(\begin{bmatrix}P&Q^{-1}&R^{-1}\end{bmatrix}\right)$
7:     $\mathbf{q}\leftarrow\begin{bmatrix}\mathbf{q}^{T}&\mathbf{0}^{T}&-(R^{-1}\mathbf{y}_{k+1})^{T}\end{bmatrix}^{T}$
@@
9:  $\mathcal{X}_{0}\leftarrow\begin{bmatrix}0&\dots&0&I\end{bmatrix}\mathcal{Z}$
@@ V-C Safety Verification
[…]
@@
[…]
@@ VI-A Implementation
The following numerical examples are implemented using the open-source toolbox ZonoOpt [^1] https://github.com/psu-PAC-Lab/ZonoOpt, which was developed as a contribution of this paper. This toolbox provides classes for zonotopes, constrained zonotopes, and hybrid zonotopes, implements the set operations in (3) (among others), and includes an implementation of Algorithm 1. ZonoOpt is implemented in C++ and Python bindings are available for all classes and functions. Sparse linear algebra is utilized throughout using the Eigen linear algebra library [39]. ZonoOpt is influenced by the MATLAB toolbox zonoLAB [1].
@@ VI-B Model Predictive Control
Consider the MPC problem described in Sec. V-A with a discrete time double integrator model
$\mathbf{u}_{k}=\begin{bmatrix}a_{x,k}&a_{y,k}\end{bmatrix}^{T}\;,$ (34b)
$A=\begin{bmatrix}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\;,\;B=\begin{bmatrix}\frac{\Delta t^{2}}{2}&0\\ 0&\frac{\Delta t^{2}}{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}\;.$ (34c)
The time-varying constraint set $\mathcal{S}_{k}$ is decomposed into a time-varying position component and a constant velocity component as $\mathcal{S}_{k}=\mathcal{P}_{k}\times\mathcal{V}$. The velocity and input constraints $\mathcal{V}$ and $\mathcal{U}$ are regular 12-sided polygons, and time-varying position constraints are denoted by $\mathcal{P}_{k}$. In this example, the $\mathcal{P}_{k}$ are given by centrally symmetric hexagons. The motivation for the regular polygon constraints comes from motion planning problems where differential flatness is used to transform the nonlinear unicycle model into the linear double integrator model [40]. In that context, it is necessary to approximate the quadratic constraints $v_{x,k}^{2}+v_{y,k}^{2}\leq v_{\text{max}}^{2}$ and $a_{x,k}^{2}+a_{y,k}^{2}\leq(v_{\text{min}}\omega_{\text{max}})^{2}$, where $v_{\text{min}}$, $v_{\text{max}}$, and $\omega_{\text{max}}$ are linear and angular velocity limits [41]. For autonomous vehicle motion planning, the time-varying position constraints $\mathcal{P}_{k}$ may be specified by a higher-level autonomous planning algorithm such as in [42].
In this example, the state vector has dimensions of $\mathbf{x}\sim\begin{bmatrix}\text{m}&\text{m}&\text{m/s}&\text{m/s}\end{bmatrix}^{T}$ and the input vector has dimensions of $\mathbf{u}\sim\begin{bmatrix}\text{m}/\text{s}^{2}&\text{m}/\text{s}^{2}\end{bmatrix}^{T}$. The cost function is defined by $Q=Q_{N}=\text{diag}\left(\begin{bmatrix}1&1&0&0\end{bmatrix}\right)$ and $R=10I$. The initial condition is $\mathbf{x}_{0}=\begin{bmatrix}0&-10&0&0\end{bmatrix}^{T}$ and the discrete time step is $\Delta t=1.0$ s. The velocity constraint set $\mathcal{V}$ is defined according to $v_{\text{max}}=5$ m/s and the input constraint set $\mathcal{U}$ is defined according to $v_{\text{min}}=0.1$ m/s and $\omega_{\text{max}}=75$ deg/s. The MPC horizon is $N=55$. In this example, $\mathcal{U}$ and $\mathcal{S}_{k}$ are centrally symmetric and as such are represented as zonotopes.
@@
$P=\text{blkdiag}\left(\begin{bmatrix}Q&R&Q&\dots&Q&R&Q_{N}\end{bmatrix}\right)\;,$ (35c)
$\mathbf{q}=\left[\begin{matrix}\mathbf{0}^{T}&\mathbf{0}^{T}&-(Q\mathbf{x}^{r}_{1})^{T}&\dots\end{matrix}\right.$
@@
Solution times for the MPC optimization problem are compared for the considered QP solvers (ZonoOpt, OSQP, and Gurobi) and problem formulations (Algorithm 2 vs. (35)) in Fig. 4. To assess the scalability of these solution approaches to large MPC horizons, the MPC problem was modified using a scaling factor $f$ such that $N=fN_{0}$ and $\Delta t=\Delta t_{0}/f$ where $N_{0}=55$ and $\Delta t_{0}=1$ s. The position constraint sets $\mathcal{P}_{k}$ and reference states $\mathbf{x}^{r}_{k}$ are correspondingly adjusted via interpolation. The MPC optimization problem was solved for $f=\{1,2,...,21\}$, resulting in MPC horizons ranging from $N=55$ to $N=1155$. Fig. 4 empirically shows that the proposed MPC problem formulation using reachability analysis of constrained zonotopes and corresponding solution approach using ADMM achieves lower solution times and improved scalability with respect to MPC horizon when copared to state-of-the-art problem formulations and QP solvers.
553306058801155Horizon $N$0050505050100100100100150150150150200200200200250250250250300300300300350350350350400400400400Solution time [ms]ZonoOpt CZOSQP CZOSQP H-repGurobi CZGurobi H-rep

### v2

@@ IV-B Infeasibility Detection
The vector $\mathbf{v}$ can be written as $\mathbf{v}=A^{T}\bm{\lambda},\;\bm{\lambda}\in^{n_{C}}$. For all $\bm{\xi}\in\mathcal{A}$, $A\bm{\xi}=\mathbf{b}$, and therefore $\mathbf{v}^{T}\bm{\xi}=\bm{\lambda}^{T}\mathbf{b}$ $\forall\bm{\xi}\in\mathcal{A}$ and $\min_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})=\max_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})=\mathbf{v}^{T}\tilde{\bm{\xi}}$. Defining $c]=\mathbf{v}^{T}[\bm{\beta}]=[\underline{c},\overline{c}]$, $\min_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})=\underline{c}$ and $\max_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})=\overline{c}$. If $\mathbf{v}^{T}\tilde{\bm{\xi}}\notin[c]$, then $\min_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})>\max_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})$ or $\min_{\bm{\xi}\in\mathcal{A}}(\mathbf{v}^{T}\bm{\xi})>\max_{\bm{\xi}\in\mathcal{B}}(\mathbf{v}^{T}\bm{\xi})$, and by the separating hyperplane theorem $\mathcal{A}\cap\mathcal{B}=\emptyset$ [40], so (23) is infeasible. ∎
@@
If (23) is infeasible and $Z^{T}\tilde{P}Z$ is positive definite where $Z$ is a basis for the null space of AA, then the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}\rightarrow\mathbf{v}$ for $k\geq k^{\prime}$ with $k^{\prime}$ sufficiently large, where $\mathbf{v}$ is an infeasibility certificate for problem (23).
@@
[…]
@@
IV-C Interpretation of ADMM Iterations
@@
IV-D ADMM Algorithm for Constrained Zonotopes
@@
[…]
@@
[…]
@@ V-A Model Predictive Control
[…]
@@
[…]
@@
[…]
@@
11:  $\mathbf{q}\leftarrow\begin{bmatrix}\mathbf{q}^{T}&\mathbf{0}^{T}&-(Q\mathbf{x}^{r}_{k})^{T}\end{bmatrix}^{T}$
@@ V-B Moving Horizon Estimation and Set-Valued State Estimation
$\mathbf{x}_{k+1}=A\mathbf{x}_{k}+B\mathbf{u}_{k}+\mathbf{w}_{k}\;,\;\mathbf{y}_{k}=C\mathbf{x}_{k}+\mathbf{v}_{k}\;,$ (28a)
$\text{s.t.}\;\forall k:\mathbf{x}_{k},\mathbf{x}_{k+1}\in\mathcal{S}\;,\;\mathbf{w}_{k}\in\mathcal{W}\;,\;\mathbf{v}_{k}\in\mathcal{V}\;,$ (28b)
subject to bounded process and measurement noise $\mathbf{w}_{k}$ and $\mathbf{v}_{k}$, respectively. The state domain set $\mathcal{S}$ is discussed in Sec. III. In the context of SVSE, constrained zonotopes have been used to find a bounding set $\mathcal{X}_{k}$ such that $\mathbf{x}_{k}\in\mathcal{X}_{k}$ [15]. While SVSE algorithms produce a set that is guaranteed to contain the state of the system, they do not in general produce a best state estimate within that set. The following presents an MHE implementation based on reachability analysis with constrained zonotopes that produces both a bounding set of system states and a best estimate of the state within that set.
@@
$\mathcal{X}_{k+1}=\Big{(}(A\mathcal{X}_{k}\oplus B\mathbf{u}_{k}\oplus\mathcal{W})\cap_{C}(\mathbf{y}_{k+1}\oplus(-\mathcal{V}))\Big{)}\cap\mathcal{S}\;,$ (29)
@@
$\mathcal{X}_{k+1}=\begin{bmatrix}0&0&I\end{bmatrix}\Big{(}(\mathcal{X}_{k}\times\mathcal{W}\times(\mathcal{S}\cap_{C}(\mathbf{y}_{k+1}\oplus(-\mathcal{V}))))\\ \cap_{\begin{bmatrix}A&I&-I\end{bmatrix}}\{-B\mathbf{u}_{k}\}\Big{)}\;.$ (30)
@@
$\mathbf{x}_{k+1}=A\mathbf{x}_{k}+B\mathbf{u}_{k}+\mathbf{w}_{k}\;,\;\mathbf{y}_{k}=C\mathbf{x}_{k}+\mathbf{v}_{k}\;,$ (31b)
@@
The process and measurement noise covariance matrices are $Q$ and $R$, respectively. The covariance of the state estimate at time step k=Nk=-N is $P_{-N}$, and the prior state estimate for k=Nk=-N is $\hat{\mathbf{x}}_{-N}$. The set of possible states at time step k=Nk=-N is $\mathcal{X}_{-N}$, which is updated recursively between MHE iterations by applying (30).
@@
[…]
@@
5:  $\begin{aligned} \mathcal{Z}\leftarrow&\left(\mathcal{Z}\times\mathcal{W}\times(\mathcal{S}\cap_{C}(\mathbf{y}_{k+1}\oplus(-\mathcal{V})))\right)\\ &\cap_{\begin{bmatrix}0&\dots&0&A&I&-I\end{bmatrix}}\{-B\mathbf{u}_{k}\}\end{aligned}$
@@
7:  $\mathbf{q}\leftarrow\begin{bmatrix}\mathbf{q}^{T}&\mathbf{0}^{T}&-(C^{T}R^{-1}\mathbf{y}_{k+1})^{T}\end{bmatrix}^{T}$
@@
[…]
@@ V-C Safety Verification
$\text{s.t.}\;\forall k:\mathbf{x}_{k},\mathbf{x}_{k+1}\in\mathcal{S}\;,\;\mathbf{w}_{k}\in\mathcal{W}\;,$ (33b)
@@
$\mathcal{X}_{k+1}=\left(\begin{bmatrix}0&0&I\end{bmatrix}(\mathcal{X}_{k}\times\mathcal{W}\times\mathcal{S})\cap_{[A_{c}\;I\;-I]}\{-B\mathbf{u}_{\mathit{ff}}\}\right)\;,$ (34)
where $A_{c}=A-BK$ and $\mathbf{u}_{\mathit{ff}}=K\mathbf{x}_{k}^{r}$. Given some unsafe set $\mathcal{O}$, possibly in a lower-dimensional space than the state vector, trajectories of the system are safe at time step $k$ if $R\mathbf{x}\notin\mathcal{O}\;\forall\mathbf{x}\in\mathcal{X}_{k}$, where $R$ maps $\mathbf{x}$ onto the safety-related subspace. This condition can be verified if Algorithm 1 finds an infeasibility certificate for (23), with $\mathcal{A}$ and $\mathcal{B}$ the equality and inequality constraints of $\mathcal{Z}=\mathcal{X}_{k}\cap_{R}\mathcal{O}$.
@@ VI-A Implementation
The following numerical examples apply the methods of Sec. V in the context of motion control, localization, and online safety verification problems for autonomous mobile robots. In all cases, the discrete-time double integrator model
$\mathbf{u}_{k}=\begin{bmatrix}a_{x,k}&a_{y,k}\end{bmatrix}^{T}\;,$ (35b)
$A=\begin{bmatrix}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\;,\;B=\begin{bmatrix}\frac{\Delta t^{2}}{2}&0\\ 0&\frac{\Delta t^{2}}{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}\;$, (35c)
is used, where $x_{k}$ and $y_{k}$ are position states, $v_{x,k}$ and $v_{y,k}$ are velocity states, and $a_{x,k}$ and $a_{y,k}$ are accelerations. Using differential flatness, the double integrator model can be shown to be equivalent to the nonlinear unicycle model, which is widely used to model autonomous mobile robots [46].
All examples are implemented using the open-source toolbox ZonoOpt [^1] https://github.com/psu-PAC-Lab/ZonoOpt, which was developed as a contribution of this paper. This toolbox provides classes for zonotopes, constrained zonotopes, and hybrid zonotopes, implements the set operations in (3) (among others), and includes an implementation of Algorithm 1. ZonoOpt is implemented in C++ and Python bindings are available for all classes and functions. Sparse linear algebra is utilized throughout using the Eigen linear algebra library [47]. ZonoOpt is influenced by the MATLAB toolbox zonoLAB [1].
@@ VI-B Model Predictive Control
Consider the MPC problem described in Sec. V-A using the discrete-time double integrator model (35). The time-varying constraint set $\mathcal{S}_{k}$ is decomposed into a time-varying position component and a constant velocity component as $\mathcal{S}_{k}=\mathcal{P}_{k}\times\mathcal{V}$. The velocity and input constraints $\mathcal{V}$ and $\mathcal{U}$ are regular 12-sided polygons, and time-varying position constraints are denoted by $\mathcal{P}_{k}$. In this example, the $\mathcal{P}_{k}$ are given by centrally symmetric hexagons. The regular polygon constraints approximate the quadratic constraints $v_{x,k}^{2}+v_{y,k}^{2}\leq v_{\text{max}}^{2}$ and $a_{x,k}^{2}+a_{y,k}^{2}\leq(v_{\text{min}}\omega_{\text{max}})^{2}$, where $v_{\text{min}}$, $v_{\text{max}}$, and $\omega_{\text{max}}$ are linear and angular velocity limits [48]. For motion planning problems, the time-varying position constraints $\mathcal{P}_{k}$ may be specified by a higher-level autonomous planning algorithm such as in [49].
[…]
@@
553306058801155Horizon NN05050100100150150200200250250300300350350Solution time [ms]ZonoOpt CZOSQP CZOSQP H-repGurobi CZGurobi H-rep
@@
[…]
$\mathbf{q}=\left[\begin{matrix}\mathbf{0}^{T}&\mathbf{0}^{T}&-(Q\mathbf{x}^{r}_{1})^{T}$$\dots\end{matrix}\right$.
@@
Solution times for the MPC optimization problem are compared for the considered QP solvers (ZonoOpt, OSQP, and Gurobi) and problem formulations (Algorithm 2 vs. (36)) in Fig. 5. To assess the scalability of these solution approaches to large MPC horizons, the MPC problem was modified using a scaling factor $f$ such that $N=fN_{0}$ and $\Delta t=\Delta t_{0}/f$ where $N_{0}=55$ and $\Delta t_{0}=1$ s. The position constraint sets $\mathcal{P}_{k}$ and reference states $\mathbf{x}^{r}_{k}$ are correspondingly adjusted via interpolation. The MPC optimization problem was solved for $f=\{1,2,...,21\}$, resulting in MPC horizons ranging from $N=55$ to $N=1155$. Fig. 5 empirically shows that the proposed MPC problem formulation using reachability analysis of constrained zonotopes and corresponding solution approach using ADMM achieves lower solution times and improved scalability with respect to MPC horizon when compared to state-of-the-art problem formulations and QP solvers.